OpenStudy (anonymous):

...Solve the following system of equations and show all work. y = 2x2 y = –3x −1

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

please help... I Dont understand

OpenStudy (welshfella):

since both expressions in x are equal to y then they are equal to each other so 2x^2 = -3x - 1

OpenStudy (welshfella):

this is a quadratic so write it as 2x^2 + 3x + 1 = 0 can you solve this for x?

OpenStudy (anonymous):

what do i do with the 2x^2

OpenStudy (welshfella):

this will factor to the form (2x + a)(x + b) = 0 to find a and b you need to find 2 numbers a and b where ab = 1 and a + 2b = 3

OpenStudy (anonymous):

okay hold on while i figure this sorry im really bad at this

OpenStudy (anonymous):

i cant im coming up with things like 2x+ 3x+ 2

OpenStudy (anonymous):

and thats not right

OpenStudy (welshfella):

what 2 numbers multiplied together give + 1? - think of the obvious ones

OpenStudy (anonymous):

-1 and -1 1 and 1

OpenStudy (welshfella):

yes either of those but also a + 2b = 3 so which ones are they?

OpenStudy (welshfella):

plug in the -1 and -1 and then the 1 and 1 and see which of these fit

OpenStudy (anonymous):

for the negatives i got -3

OpenStudy (anonymous):

and the positive i got 2

OpenStudy (welshfella):

2 for the positive? try again

OpenStudy (welshfella):

you need it to be = 3

OpenStudy (anonymous):

yes now ive got it lol i must of did it wrong

OpenStudy (anonymous):

now that we have that info what do i do next

OpenStudy (welshfella):

you now have the values of a and b so we can factor the equation (2x + 1)(x + 1) = 0 now when 2 things are multiplied and the result is 0 then either one of these can be 0 so we have 2x + 1 = 0 or x + 1 = 0 so you can solve these 2 equations to give 2 values of x then using these values of x plug them into one of the original equations to give the corresponding values of y

OpenStudy (anonymous):

okay thank you .